OpenStudy (anonymous):

Evaluate the function H(x) = 18 – x, when x = –6 A 108 B 12 C –24 D 24

OpenStudy (anonymous):

H(x)=18-(-6) =18+6 =24

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D lol